OpenStudy (anonymous):

Suzanne finances $315,000 with a 15/4 balloon mortgage at 3.75%. How much will she pay for principal and interest over the life of the loan?

OpenStudy (anonymous):

First find the monthly payment over 15 years ... P = 2,290.75/mo Next, find the outstanding balance after the 4 x 12 = 48th payment. Bal = 355,125.64 (this is the balloon payment) Total principal and interest = 355,125.64 + 2,290.75(48) = 357,416.36

OpenStudy (anonymous):

@Chooch146 thank you so much!

OpenStudy (anonymous):

your welcome!!!! medal??

OpenStudy (anonymous):

Thanks:]

OpenStudy (anonymous):

no thank you :) @Chooch146

OpenStudy (anonymous):

:P